OpenStudy (anonymous):

The expression (secx + tanx)2 is the same as _____.

OpenStudy (imstuck):

Since secx is the same as 1/cosx, and tanx is the same as sinx/cosx, we have 1/cosx, sinx/cosx. We can add the numerators now since the denominator is the same and we get (1+sinx)/cosx. What is that 2 stuck at the end? Does it get distributed through or something? Either way that is an equivalent expression to your given.

OpenStudy (jdoe0001):

hmmm have you covered the binomial theorem? I assume you should have by now

OpenStudy (jdoe0001):

\(\large (a\pm b)^2\implies a^2\pm 2ab+b^2\)

OpenStudy (jdoe0001):

\(\bf \large [{\color{brown}{ sec(x)}}+{\color{blue}{ tan(x)}}]^2\implies {\color{brown}{ \square }}^2+2{\color{brown}{ \square }}{\color{blue}{ \square }}+{\color{blue}{ \square }}^2\)

OpenStudy (jdoe0001):

check your choices, it'd be there

OpenStudy (anonymous):

sec^2x+2cscx+tan^2 is that the answer

OpenStudy (jdoe0001):

cscx?

OpenStudy (jdoe0001):

hmmm

OpenStudy (anonymous):

secx=1/cosx tanx=sinx/cosx so: \(=(\huge\frac{ 1+sinx}{cosx})^2\) now this you can put in many ways: for example using \((a+b)^2=a^2+2ab+b^2\) you can get \((1+2sinx+sin^2x)/(cos^2x)\) which is your answer wit coscx changed to secx

OpenStudy (anonymous):

whats the answer

OpenStudy (jdoe0001):

well. if you expand that first, waht would you get?